# Optimal Trudinger-Moser inequalities on complete noncompact Riemannian manifolds: Revisit of the argument from the local inequalities to global ones

### v1

@@ Abstract.
The main purpose of this short note, on the one hand, to is clarify some part of the proof of Theorem 1.3 in [8] in a simple way, and on the other hand, to give an alternative argument from local inequalities to global ones.
@@ 1. Introduction
This sort of ideas has also been used in our proofs of the critical Trudinger-Moser inequality (Theorem 1.3) on complete and noncompact Riemannian manifolds [8]. In the proof of Theorem 1.3 in [8], some clarifications need to be made about the uniformity of the upper bound of the Trudinger-Moser inequality in terms of the bounded domains of the level sets involved. To avoid the complexity, we will then simply give another simple proof by only establishing the uniformity of the bounds of the Trudinger-Moser inequality on geodesic balls within the manifolds (still going through the same proof of Theorem 3.1 in [8] on the geodesic balls or simply applying Theorem 3.1 in [8] on geodesic balls where the Green functions have uniform estimates) together with a Gromov’s covering lemma (similar to the proof of subcritical case in [8]). The proof of Theorem 3.1 in [8] remains valid for geodesic balls with uniform constants for the Trudinger-Moser inequalities for such balls independent of the location of the balls. Since the main purpose of this short note is to clarify the proof of Theorem 1.3 in [8] and in the meantime to give an alternative proof from the level set argument passing from the local inequalities to global one, we have chosen not to give a comprehensive account of the subject (see e.g. [8] for relevant references in the literature).
2. Some useful known lemmas
We recall the critical Moser-Trudinger inequality in [8].
@@ 2. Some useful known lemmas
The Ricci lower bound condition guarantees a finite overlapping covering of $M$. This is the so-called Gromov’s covering lemma, which states as follows:
Lemma 2.1 (Hebey [4], Lemma 1.6).
Let $(M,g)$ be a complete Riemannian manifold with Ricci lower bound $\lambda$, let $\rho>0$ be given, then there exists a sequence of points $\{x_{j}\}$ on $M$ such that for any $r\geq\rho$:
@@
    the family $\{B_{r}(x_{j})\}$ is a uniformly locally finite covering of $M$, there exists an upper bound for the overlapping number in terms of $n,\rho,r,\lambda$;
@@
On the other hand, our geometric assumptions guarantee the existence of harmonic coordinates whose harmonic radius is uniformly bounded from below. To be precise, we have
Lemma 2.2 (Hebey [4], Theorem 1.3).
Let $\alpha\in(0,1)$, $Q>1$, $\delta>0$ and $(M,g)$ satisfy our assumptions. $\Omega(\delta)=\{x\in M:\textit{dist}(x,\Omega)<\delta\}$. Then there exists a positive constant $C=C(n,Q,\alpha,\delta,i,\lambda)$ such that for any $x\in\Omega$, the $C^{0,\alpha}$ harmonic radius $r_{H}(Q,0,\alpha)(x)\geq C$.
@@
3. Proof of Theorem 2.1
To give another proof of Theorem 2.1, we achieve this through 2 steps: establishing the uniform Trudinger-Moser inequality on each $B_{\delta}(x_{j})$; gluing all inequalities from Step 1 so we extend our inequality from the local inequalities to a global one. We first prove
Theorem 3.1.
@@
The proof of Theorem 3.1 consists of 2 steps:
Step 1: Since $u\in W^{1,n}_{0}(B_{\delta}(x_{j}))$, we recall the pointwise representation formula of $u(x)$ in terms of the Dirichlet Green’s function:
@@
The existence of the Green’s function comes from a classical paramatrix argument (see Aubin, Nonlinear Analysis on Manifolds. Monge-Ampére Equations, 1982, Chapter 4). To be precise, for some positive integer $k>\frac{n}{2}$,
@@
where $C=C(n,\lambda,Q,\delta)$. Indeed, from the definition of $\Gamma_{l}(x,y)$, their singularity decrease as $l$ grows hence we first look at
@@
For the first term of the right hand side,
@@
$\leq$ $Q^{n/2}\omega_{n-1}\rho^{\beta-n}\int_{0}^{\rho}r^{\alpha-n}r^{n-1}dr,$
$\leq$ $Q^{n/2}\omega_{n-1}\rho^{\alpha+\beta-n},$
where in the third line we use the property of harmonic coordinates $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$. Similarly we can estimate the second term of the right hand side of (3.3). For the third term, we have
@@
(3.7) $\leq$ $C_{n}Q^{n/2}\omega_{n-1}\int_{3\rho}^{\delta}r^{\alpha+\beta-2n}r^{n-1}dr,$
(3.8) $\leq$ $C_{n,\alpha,\beta}Q^{n/2}\omega_{n-1}((3\rho)^{\alpha+\beta-n}-\delta^{\alpha+\beta-n}),$
(3.9) $\leq$ $C_{n,\alpha,\beta}Q^{n/2}\omega_{n-1}\rho^{\alpha+\beta-n}.$
From Lemma 3.1, $\Gamma_{k+1}$ together with its derivatives, are uniformly bounded by a constant $C=C(n,\lambda,i,Q)$. Therefore, to obtain the uniform estimate for Green’s function, we need to prove the uniform $C^{1}$ estimate of $F(x,y)$.
@@
Therefore we obtain the $C^{0}$ estimate of $F$. Now we recall the Schauder interior estimate of elliptic equations (see Gilbarg-Trudinger [3], Theorem 6.2):
Theorem 3.2 (Schauder estimates).
Let $u\in C^{2,\alpha}(\Omega)$ be a bounded solution in $\Omega$ and the equation
$a^{ij}\partial_{ij}u+b^{i}\partial_{i}u+cu=f,$
where $f\in C^{\alpha}$ and there are positive constants $C_{1},C_{2}$ such that $a^{ij}\xi_{i}\xi_{j}\geq C_{1}|\xi|^{2}$ and
$|a^{ij}|_{C^{0,\alpha}},\ |b^{i}|^{(1)}_{C^{0,\alpha}},\ |c|^{(2)}_{C^{0,\alpha}}\leq C_{2}.$
@@
In our case, $b^{i},c$ vanish and from Theorem 2.2 in harmonic coordinates $a^{ij}=g^{ij}$ belongs to $C^{0,\alpha}$. $f$ corresponds to $\Gamma_{k}$, which is proved to be uniformly bounded from Lemma 3.1. Therefore we have $|F|^{*}_{C^{2,\alpha}}\leq C(n,\lambda,i,Q,\alpha)$. In specific, $\nabla_{y}F(x,y)||_{\infty}\leq C(n,\lambda,i,Q)$ and we complete the proof. ∎
Step 2: We use Adams’ method to obtain the equivalent version of Theorem 3.1:
@@
The proof of Theorem 3.3 is exactly the same as that of Theorem 3.1 in [8] which is one of the main results. The parameters $\lambda,i$ do not explicitly appear in the estimates, but will determine $Q$ and $r_{H}(\alpha,0,Q)$, which gives the bounds of $\delta$. This shows the constants in the Moser-Trudinger inequality on any geodesic balls only depend on the radius $\delta$ which in turn depends on $\lambda$ and $i$.

### v2

@@ Abstract.
The main purpose of this short note, on the one hand, is to clarify some part of the proof of Theorem 1.3 in [10] using the level set argument in a simple way, and on the other hand, to give an alternative argument from local inequalities to global ones by using Gromov’s covering lemma. Moreover, we show that the Green functions for the Dirichlet problem on any small geodesic balls (or harmonic coordinate charts) on Riemannian manifolds with positive injective radius and Ricci lower bound have uniform estimates independent of the locations of the balls (or the charts). This will allow us to conclude that the Trudinger-Moser inequality on any level set is uniformly bounded independent of the location of the level set and thus rigorize the proof of Theorem 1.3 in [10]. This latter result of Green’s function estimates is of its independent interest and can be found useful elsewhere.
@@ 1. Introduction
This sort of ideas has also been used in our proofs of the critical Trudinger-Moser inequality (Theorem 1.3) on complete and noncompact Riemannian manifolds [10]. To be precise, we established
@@
In the proof of Theorem 1.1 some clarifications need to be made about the uniformity of the upper bound of the Trudinger-Moser inequality in terms of the bounded domains of the level sets involved. In this note, we will then simply give another simple proof by only establishing the uniformity of the bounds of the Trudinger-Moser inequality on geodesic balls within the manifolds (still going through the same proof of Theorem 3.1 in [10] on the geodesic balls or simply applying Theorem 3.1 in [10] on geodesic balls where the Green functions have uniform estimates) together with a Gromov’s covering lemma (similar to the proof of subcritical case in [10]). The proof of Theorem 3.1 in [10] remains valid for geodesic balls with uniform constants for the Trudinger-Moser inequalities for such balls independent of the location of the balls.
The main result of this note is that the Green functions for the Dirichlet problem on any small geodesic balls on Riemannian manifolds with positive injective radius and Ricci lower bound have uniform estimates independent of the locations of the balls. This will allow us to conclude that the Trudinger-Moser inequality on any level set is uniformly bounded independent of the location of the level set. More precisely, in the proof of Theorem 1.3 in [10], the level set $\Omega(u)$ can be covered by a finite number of small geodesic balls (or harmonic coordinate charts) and the number of such balls (or charts) is uniformly bounded. Therefore, the same level argument from local inequalities to the global ones in [10] are justified.
Since the main purpose of this short note is to rigorize the proof of Theorem 1.3 in [10] and in the meantime to give an alternative proof from the level set argument passing from the local inequalities to global one, we have chosen not to give a comprehensive account of the subject (see e.g. [10] for relevant references in the literature).
2. Some preliminaries in differential geometry
We first recall the Gromov’s covering lemma:
Lemma 2.1 (Lemma 1.6 of Hebey [6]).
Let $(M,g)$ be a complete Riemannian manifold with Ricci lower bound $\lambda$, let $\delta>0$ be given, then there exists a sequence of points $\{x_{j}\}$ on $M$ such that for any $r\geq\delta$:
@@ 2. Some preliminaries in differential geometry
    the family $\{B_{r}(x_{j})\}$ is a uniformly locally finite covering of $M$, there exists an upper bound for the overlapping number in terms of $n,\delta,r,\lambda$;
@@
On the other hand, lower bounds for both Ricci curvature and injectivity radius guarantee the existence of $C^{0,\alpha}$ harmonic coordinates. This is due to Anderson-Cheeger [1] (see also Hebey [6]) and to be precise:
Lemma 2.2.
Let $\alpha\in(0,1)$, $Q>1$, $\delta>0$ and $(M,g)$ satisfy $\textit{Ric}\geq\lambda g$ and $\textit{inj}(M,g)\geq i>0$. Then there exists a positive constant $C=C(n,Q,\alpha,\delta,i,\lambda)$ such that for any $x\in\Omega$, the $C^{0,\alpha}$ harmonic radius $r_{H}(Q,0,\alpha)(x)\geq C$.
@@
  • In the overlapping charts $B_{\delta}(x_{i})\cap B_{\delta}(x_{j})$, the transition function satifies
    $y^{k}\circ(x^{l})^{-1}||_{C^{1,\alpha}}\leq C(n,Q,\lambda,\alpha,\delta,i).$
@@
Remark 2.1.
Lemma 2.2 is due to Anderson-Cheeger [1]. The $C^{0,\alpha}$ result was achieved by the existence of $W^{1,p}$- harmonic coordinates and a classical Sobolev embedding for $\alpha=1-\frac{n}{p}$. This means in each local chart, the metric tensor satisfies
@@
Similar embedding also appears for the coordinate functions $x^{i},i=1,\dots,n$ and we have $W^{2,p}$ uniform bounds.
3. Green’s function estimate
For $u\in W^{1,n}_{0}(B_{\delta}(x_{j}))$, we recall the pointwise representation formula of $u(x)$ in terms of the Dirichlet Green’s function:
@@ 3. Green’s function estimate
The existence of the Green’s function comes from a classical paramatrix argument (see e.g. Chapter 4 of Aubin [2]). To be precise, for some positive integer $k>\frac{n}{2}$,
@@
where $C=C(n,\lambda,Q,\lambda,\delta)$. Indeed, from the definition of $\Gamma_{l}(x,y)$, their singularity decrease as $l$ grows hence we first look at
@@
It is well known that in polar coordinates
$dV=r^{n-1}\left(1-\frac{\textit{Ric}(\theta)}{3}r^{2}+o(r^{2})\right)d\theta dr.$
When $\delta$ is small, providing Ricci curvature lower bounded, the first term of the right hand side,
@@
$\leq$ $C(\lambda,\delta)\omega_{n-1}\rho^{\beta-n}\int_{0}^{\rho}r^{\alpha-n}r^{n-1}dr,$
$\leq$ $C(\lambda,\delta)\omega_{n-1}\rho^{\alpha+\beta-n},$
Similarly we can estimate the second term of the right hand side of (3.3). For the third term, we have
@@
(3.7) $\leq$ $C_{n}C(\lambda,\delta)\omega_{n-1}\int_{3\rho}^{\delta}r^{\alpha+\beta-2n}r^{n-1}dr,$
(3.8) $\leq$ $C_{n,\alpha,\beta,\lambda,\delta}\omega_{n-1}((3\rho)^{\alpha+\beta-n}-\delta^{\alpha+\beta-n}),$
(3.9) $\leq$ $C_{n,\alpha,\beta,\lambda,\delta}\omega_{n-1}\rho^{\alpha+\beta-n}.$
It is easy to verify that in polar coordinates,
[…]
where $r=d(x,y)$ and $|\widetilde{g}|=\textit{det}(\widetilde{g}_{ij})$ denotes the determinant of the corresponding metric tensor. By carefully choosing the smooth cut-off function $f$, both the first two terms $|(n-3)f^{\prime}-rf^{\prime\prime}|$ and the coefficient of the third term $((n-2)f-rf^{\prime})$ can be uniformly bounded by $C(n,\delta,i)$. With the help of Lemma 3.1, we only need to consider the integral of the form
$\int_{B_{\delta}(x_{j})}|\nabla_{z}\ln\sqrt{|\widetilde{g}|}|(z)H(z,y)dV_{z}.$
An easy change of variablse from polar coordinates to harmonic coordinates shows
$\widetilde{g}_{ij}=g_{kl}\frac{\partial x^{k}}{\partial\widetilde{x}^{i}}\frac{\partial x^{l}}{\partial\widetilde{x}^{j}},$
where $\{x^{k}\}$ and $\{\widetilde{x}^{i}\}$ denotes the coordinate variables respectively. For simplicity we denote the transition matrix $A=\left(\frac{\partial x^{k}}{\partial\widetilde{x}^{i}}\right)$ then
$|\nabla\ln\sqrt{|\widetilde{g}|}|\leq\frac{|\nabla|g||}{2|g|}+\frac{|\nabla|A||}{|A|}.$
Since in harmonic coordinates $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$, we have
$\left|\int_{B_{\delta}(x_{j})}\frac{|\nabla|g||}{|g|}(z)H(z,y)dV_{z}\right|$ $\leq C(n,Q)\int_{B_{\delta}(x_{j})}\frac{|\nabla g(z)|}{d(z,y)^{n-2}}dV_{z}$
Similar to the argument in the proof of Lemma 3.1, the integral can be splitted into 3 pieces. Here we just check the part over $B_{\rho}(x)\cap B_{\delta}(x_{j})$, the other two parts can be treated in a similar manner:
$\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}\frac{|\nabla g(z)|}{d(z,y)^{n-2}}dV_{z}$ $\leq\frac{C(n)}{d(x,y)^{n-2}}\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}|\nabla g|dV_{z};$
[…]
$\leq\frac{C(n,Q)}{d(x,y)^{n-2-\frac{n}{p^{\prime}}}}.$
On the other hand, $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$ implies there exists $C(n,Q)>0$ such that $|A|\geq C(n,Q)>0$. Hence we have
$\left|\int_{B_{\delta}(x_{j})}\frac{|\nabla|A||}{|A|}(z)H(z,y)dV_{z}\right|\leq C(n,Q)\int_{B_{\delta}(x_{j})}\frac{|\nabla|A|(z)|}{d(z,y)^{n-2}}dV_{z}$
We can estimate the above integral in a similar manner, but this time we make use of the $W^{2,p}$ estimate of the harmonic coordinate function $x^{i},i=1,\dots,n$. Indeed, in the derivation of the existence of harmonic coordinates, the $L^{p}$ theory of elliptic operators gives the uniform bound of $x^{i}||_{W^{2,p}}$, where $p$ satisfies $\alpha=1-\frac{n}{p}$. Hence there exists $C(n,\lambda,i,Q)$ such that $||\nabla A|_{L^{p}}\leq C(n,\lambda,i,Q)$.
Therefore from Lemma 3.1, $\Gamma_{k+1}$ together with its derivatives, are uniformly bounded by a constant $C=C(n,\lambda,i,Q)$. To obtain the uniform estimate for Green’s function, it suffices to prove the uniform $C^{1}$ estimate of $F(x,y)$.
@@
Therefore we obtain the $C^{0}$ estimate of $F$. Now we proceed the $C^{1,\alpha}$ Schauder estimate on $B_{\delta}(x_{j})$ so we obtain the $L^{\infty}$ estimate of $\nabla_{y}F(x,y)$. The proof reduces the problem to Euclidean boundary $C^{1,\alpha}$ estimates via harmonic coordinates and a flattening of the boundary. The argument is classical and standard, we only sketch the proof.
Step 1: Reduction to a Euclidean domain. Let $\Phi:B_{r}(p)\to\Omega\subset\mathbb{R}^{n}$ be the harmonic coordinate chart. Let $\widetilde{u}=u\circ\Phi^{-1}$, $\widetilde{f}=f\circ\Phi^{-1}$, and $\widetilde{\varphi}=\varphi\circ\Phi^{-1}$. Then $\widetilde{u}$ solves a uniformly elliptic equation in divergence form on $\Omega$:
@@
with boundary data $\widetilde{u}=\widetilde{\varphi}$ on $\partial\Omega$. The coefficients satisfy
$\widetilde{a}^{ij},\widetilde{b}^{i},\widetilde{c}\in C^{\alpha}(\overline{\Omega}),\qquad\|\widetilde{a}^{ij}\|_{C^{\alpha}}+\|\widetilde{b}^{i}\|_{C^{\alpha}}+\|\widetilde{c}\|_{C^{\alpha}}\leq C(n,K,i_{0},r).$
Step 2: Flattening the boundary. Since $r<i_{0}$, the boundary $\partial B_{r}(p)$ is smooth, and thus $\partial\Omega$ is $C^{1,\alpha}$. Cover $\partial\Omega$ by finitely many coordinate patches $\{U_{k}\}$ such that each patch is mapped by a $C^{1,\alpha}$-diffeomorphism $\Psi_{k}$ to a half-ball $B_{1}^{+}\subset\mathbb{R}^{n}$. Let $\{\eta_{k}\}$ be a partition of unity subordinate to this cover.
Step 3: Apply Euclidean boundary $C^{1,\alpha}$ estimates. Fix one patch $U_{k}$ and define $v_{k}=\widetilde{u}\circ\Psi_{k}^{-1}$ on $B_{1}^{+}$. Then $v_{k}$ solves a uniformly elliptic equation in divergence form with $C^{\alpha}$ coefficients and boundary data $v_{k}=\varphi_{k}$ on $\partial B_{1}^{+}$. By the Euclidean boundary $C^{1,\alpha}$ estimate (Gilbarg–Trudinger [5]), we have
$\|v_{k}\|_{C^{1,\alpha}(\overline{B_{1/2}^{+}})}\leq C\Big(\|v_{k}\|_{L^{\infty}(B_{1}^{+})}+\|\widetilde{f}\|_{C^{\alpha}(B_{1}^{+})}+\|\varphi_{k}\|_{C^{1,\alpha}(\partial B_{1}^{+})}\Big).$
Step 4: Patch local estimates to the whole domain. Multiplying by $\eta_{k}$ and summing over $k$, we obtain a global estimate on $\Omega$:
$\|\widetilde{u}\|_{C^{1,\alpha}(\overline{\Omega})}\leq C\Big(\|\widetilde{u}\|_{L^{\infty}(\Omega)}+\|\widetilde{f}\|_{C^{\alpha}(\Omega)}+\|\widetilde{\varphi}\|_{C^{1,\alpha}(\partial\Omega)}\Big).$
Using the uniform $C^{1,\alpha}$ control of the coordinate chart $\Phi$, we obtain the desired estimate.
4. Proof of Theorem 1.1
To prove Theorem 1.1, we achieve this through 2 steps: establishing the uniform Trudinger-Moser inequality on each $B_{\delta}(x_{j})$; gluing all inequalities from Step 1 so we extend our inequality from the local inequalities to a global one. We first prove
Theorem 4.1.
[…]
We use Adams’ method to obtain an equivalent version of Theorem 4.1:
@@
With the help of Green’s function estimates in the last section, the proof of Theorem 4.2 is exactly the same as that of Theorem 3.1 in [10] which is one of the main results. The parameters $\lambda,i$ do not explicitly appear in the estimates, but will determine $Q$ and $r_{H}(\alpha,0,Q)$, which gives the bounds of $\delta$. This shows the constants in the Moser-Trudinger inequality on any geodesic balls only depend on the radius $\delta$ which in turn depends on $\lambda$ and $i$.
@@
Here we make some remarks about our alternative proof.
Remark 4.1.
To establish global inequalities from local ones, we choose appropriate cutoff functions $\psi_{j}$ such that $\nabla(\psi_{j}^{2}u)||_{L^{n}}\leq 1$. Thus the scale parameter $\delta$ actually depends on $\tau$. On the other hand, in the Green’s function estimates, we make use of both $W^{1,p}$ bounds of $g$ and $W^{2,p}$ bounds of $x^{i}$. There we need to choose proper value of $p$, hence $\alpha$. This can be done since the harmonic coordinates exists for arbitrary chosen $\delta$’s and $\alpha$’s. In other words, our inequalities can only be proved in specific harmonic coordinates.
Remark 4.2.
There is another version of harmonic coordinates one can use to prove our main theorem. In [1], it reads as a collection of finitely overlapping charts $\{U_{j},F_{j}\}$, where $F_{j}:U_{j}\to\mathbb{R}^{n}$ is of the form $U_{j}=F_{j}^{-1}(B_{r_{H}}(0))$. In such choice of harmonic coordinates, the a priori estimate of $F(x,y)$ will be significantly simplified since the Dirichlet problem is taken on Euclidean balls whose boundary is smooth and can be uniformly finitely covered. On the other hand, the domains $\{U_{j}\}$ share the same Gromov’s covering properties. Noticing that $\textit{Vol}(U_{j})=\int_{U_{j}}dV=\int_{B_{r_{H}}(0)}\sqrt{|g|}dx\sim C(n,Q)r_{H}^{n}$, it is still possible to perform the local to global argument using this collection of charts, instead of geodesic balls.